OpenStudy (anonymous):

I have to solve for x but when I try I get confused and get it wrong. HELP PLZ

OpenStudy (misty1212):

HI!

OpenStudy (anonymous):

plz wait, I have to post it

OpenStudy (misty1212):

k dear i will wait

OpenStudy (anonymous):

x+(2x-3)greater than or equal to 4

OpenStudy (misty1212):

\[x+(2x-3)\geq 4\]

OpenStudy (anonymous):

yes that's the equation ^

OpenStudy (misty1212):

ok we can combine like terms first \[3x-3\geq 4\]

OpenStudy (misty1212):

add \(3\) \[3x\geq 7\]

OpenStudy (anonymous):

ohh.

OpenStudy (misty1212):

divide by 3 \[x\geq \frac{7}{3}\]

OpenStudy (misty1212):

you have a cute picture \(\color\magenta\heartsuit\)\(\color\magenta\heartsuit\)

OpenStudy (anonymous):

Thanks, I drew it

OpenStudy (misty1212):

i wish i could draw

OpenStudy (anonymous):

XD also one question, do I leave it 7 over 3 or divide 7 by 3?

OpenStudy (misty1212):

i would leave it

OpenStudy (anonymous):

ok, got it and I wish I was as good at math as in drawing

OpenStudy (anonymous):

Thank You for the help!! =)

OpenStudy (misty1212):

\(\color\magenta\heartsuit\)\(\color\magenta\heartsuit\)